OpenStudy (anonymous):

Calculate the mass in grams of 3.82 mol of (NH4)2SO3

OpenStudy (anonymous):

n= m/M -> m = n * M use periodic table of elements to get M and just put numbers in that eqation and you are done

OpenStudy (anonymous):

mass of 1mole of (NH4)2SO3 is 116g mass of 3.82 moles is 116*3.82 = 443.12g